OpenStudy (jkbo):

Please help use the quadratic formula to solve the equation 4x^2+6x-4=0

OpenStudy (anonymous):

do you know what the quadratic formula is?

OpenStudy (jkbo):

No I just learned this

OpenStudy (anonymous):

ok so the quadratic formula is \[\frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

\[ax ^{2}+b x+c=0\] Quadratic formula \[x=\frac{ -b \pm \sqrt{ b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

now that you have that can you tell me what a=? b=? and c=?

OpenStudy (jkbo):

4,6,4

OpenStudy (anonymous):

close but its actually 4,6,(-)4

OpenStudy (anonymous):

a=4,b=6,c=-4 plug in the values and find the solution

OpenStudy (anonymous):

now that you have those you just plug and chug

OpenStudy (jkbo):

I got -3.25

OpenStudy (jkbo):

@nickolgee429

OpenStudy (anonymous):

you should get 2 answers because it is b +- the square root

OpenStudy (anonymous):